For a confidence level of 96% with a sample size of 55, find the z score for the right bound of the confidence interval. This requires taking NORMINV of the appropriate value. Show your answer to 3 decimals.
Answer #1

Solution:

We are given

Confidence level = 96%

α = 1 - 0.96 = 0.04

α/2 = 0.04/2 = 0.02

For right side or upper 0.02 area, the critical Z value is given as below:

Z = 2.053749

Answer: Z = 2.054